OpenStudy (anonymous):

The line containing the diagonal of a square whose vertices are A(-3, 3), B(3, 3), C(3, -3), and D(-3, -3). Find two equations, one for each diagonal. please help me :(*

OpenStudy (anonymous):

 \[Eq.of AC is y-y1=\frac{ y2-y1 }{x2-x1 } \left( x-x1 \right)\] similarly eq. of BD

OpenStudy (anonymous):